OpenStudy (anonymous):

I'm having trouble with simple arithmetic : \[c_1+c_2=\frac{13}{10}\] \[-2c_1+c_2=\frac{8}{10}\]

OpenStudy (n00bstyle):

Always think of substituting the one into the other. First of all make sure to have as many unknows as you have eqations!

OpenStudy (n00bstyle):

In this case you have two unknows (c1 and c2), and two equations. Calculate c1 from the first equation for example (c1 = 13/10 - c2). Plug this into equation 2 and you can calculate c2. After this, when c2 is known, you can plug it back into c1 = 13/10 - c2

hartnn (hartnn):

subtract the equations, whats the problem ??

OpenStudy (anonymous):

I know how to do substitution...my solution is different from chegg.com....that's why I want to see what other people are getting

hartnn (hartnn):

c1 =1/6 ??

OpenStudy (n00bstyle):

and c2 is 17/15 I guess...

hartnn (hartnn):

the Q u asked and the pic have different equations...

OpenStudy (anonymous):

sorry when I did the auxiliary equation, initially, I had c_1 and they called it c_2 and vice versa

hartnn (hartnn):

so, c1 = 1/6 and c2=17/15 for the Q u posted.

OpenStudy (n00bstyle):

You made a mistake: add -1/2 - 3/10 = -5/10 - 3/10 = -8/10 In my opinion...

OpenStudy (anonymous):

I got 1/6 finally!!!!!!

OpenStudy (anonymous):

thanks guys, for a second I thought that the person posting on chegg.com made a mistake...turns out it was my mistake

OpenStudy (n00bstyle):

Well...it isn't rocketscience :D Alright, good ;)